### v1

@@ Abstract
The increasing complexity and cost of clinical trials, particularly in the context of oncology and advanced therapies, pose significant challenges for drug development. This study evaluates the predictive capabilities of large language models (LLMs) such as GPT-3.5, GPT-4, and HINT in determining clinical trial outcomes. By leveraging a curated dataset of trials from ClinicalTrials.gov, we compare the models’ performance using metrics including balanced accuracy, specificity, recall, and Matthews Correlation Coefficient (MCC). Results indicate that GPT-4o demonstrates robust performance in early trial phases, achieving high recall but facing limitations in specificity. Conversely, the HINT model excels in recognizing negative outcomes, particularly in later trial phases, offering a balanced approach across diverse endpoints. Oncology trials, characterized by high complexity, remain challenging for all models. Additionally, trial duration and disease categories influence predictive performance, with longer durations and complex diseases such as neoplasms reducing accuracy. This study highlights the complementary strengths of LLMs and HINT, providing insights into optimizing predictive tools for clinical trial design and risk management. Future advancements in LLMs are essential to address current gaps in handling negative outcomes and complex domains.
@@ 2.1 Prediction on the Entire Clinical Trial Dataset
The test results across all models on the complete dataset are shown in the table below. We observed that the GPT-4o model performed the best in terms of balanced accuracy and MCC, reaching a balanced accuracy of 0.573 and an MCC of 0.212(Table 1). Although its recall was high at 0.931, its specificity remained relatively low at 0.214, suggesting a tendency to over-classify trials as successful while overlooking some potential failed trials.
GPT-4mini and GPT-4, while achieving perfect recall (1), exhibited extremely low specificity (0 and 0.083, respectively), indicating a strong bias toward predicting all cases as positive, resulting in balanced accuracies of 0.500 and 0.542. GPT-3.5 also demonstrated this trend with a recall of 0.997 and a specificity of 0.011, leading to a balanced accuracy of 0.504.
The llama3 Llama3 model achieved moderate recall (0.949) but low specificity (0.085), resulting in a balanced accuracy of 0.517 and a low MCC of 0.058, indicating limited effectiveness in correctly identifying negative samples.
Among all models, the HINT model achieved the highest specificity at 0.541, which contributed to its relatively strong balanced accuracy of 0.563 and an MCC of 0.111. Despite having lower recall and balanced accuracy compared to other models, HINT’s ability to identify negative cases suggests a more balanced performance across positive and negative samples.
@@
| gpt4mini | 0.500 | 0.000 | 1.000 | 0.000 |
| gpt4o | 0.573 | 0.212 | 0.931 | 0.214 |
| gpt4 | 0.542 | 0.234 | 1.000 | 0.083 |
| gpt3.5 | 0.504 | 0.049 | 0.997 | 0.011 |
| HINT | 0.563 | 0.111 | 0.586 | 0.541 |
| llama3 | 0.517 | 0.058 | 0.949 | 0.085 |
@@ 2.2 Prediction of Different Clinical Trial Phases
As clinical trials progress from Phase I to Phase III, different models exhibit varying trends in metrics such as balanced accuracy, Matthews correlation coefficient (MCC), and specificity. GPT-4o demonstrates relatively stable balanced accuracy across all phases, while the HINT model shows a gradual improvement in specificity.
In Phase I (Table 2), GPT-4o and GPT-4 achieve relatively high balanced accuracies of 0.557 and 0.600, respectively, but their specificities are low, at 0.40 and 0.20. This indicates certain limitations in identifying negative samples. GPT-3.5 and llama3 also have low specificities, and GPT-4mini has a specificity of 0, showing a complete lack of ability to recognize negative samples. In contrast, the HINT model attains a balanced accuracy of 0.562 and a specificity of 0.50, demonstrating certain advantages in balancing the detection of positive and negative samples.
In Phase II, GPT-4o maintains its balanced accuracy at 0.556, exhibiting stability, and its MCC increases to 0.277, although its specificity only slightly rises to 0.111. GPT-4’s balanced accuracy slightly improves to 0.571, with an MCC of 0.319 and a marginal increase in specificity to 0.143. The HINT model’s balanced accuracy slightly decreases to 0.516, but its specificity further increases to 0.532, indicating continuous improvement in identifying negative samples. GPT-3.5 and llama3 show minimal gains in specificity, and their ability to recognize negative samples remains limited.
By Phase III, GPT-4o achieves its best performance, with balanced accuracy increasing to 0.667 and MCC reaching 0.509. Its specificity also improves to 0.333, continuing to display relatively stable performance. The HINT model further enhances its performance in this phase, achieving a balanced accuracy of 0.699 and a significant increase in specificity to 0.696. This demonstrates a balanced capability in recognizing both positive and negative samples. GPT-3.5 and llama3 still exhibit low specificities and weak abilities in identifying negative samples, while GPT-4mini’s specificity remains at 0 across all phases.
@@
| 1 | gpt4mini | 0.500 | 0.000 | 1.000 | 0.000 |
| 2 | gpt4o | 0.557 | 0.120 | 0.714 | 0.400 |
| 3 | gpt4 | 0.600 | 0.333 | 1.000 | 0.200 |
| 4 | gpt3.5 | 0.498 | -0.034 | 0.997 | 0.000 |
| 5 | HINT | 0.562 | 0.126 | 0.625 | 0.500 |
| 6 | llama3 | 0.538 | 0.121 | 0.927 | 0.150 |
@@
| 1 | gpt4mini | 0.500 | 0.000 | 1.000 | 0.000 |
| 2 | gpt4o | 0.556 | 0.277 | 1.000 | 0.111 |
| 3 | gpt4 | 0.571 | 0.319 | 1.000 | 0.143 |
| 4 | gpt3.5 | 0.505 | 0.060 | 0.997 | 0.013 |
| 5 | HINT | 0.516 | 0.030 | 0.500 | 0.532 |
| 6 | llama3 | 0.526 | 0.086 | 0.943 | 0.109 |
@@
| 1 | gpt4mini | 0.500 | 0.000 | 1.000 | 0.000 |
| 2 | gpt4o | 0.667 | 0.509 | 1.000 | 0.333 |
| 3 | gpt4 | 0.622 | 0.394 | 0.995 | 0.250 |
| 4 | gpt3.5 | 0.502 | 0.023 | 0.995 | 0.009 |
| 5 | HINT | 0.699 | 0.312 | 0.703 | 0.696 |
| 6 | llama3 | 0.512 | 0.048 | 0.966 | 0.059 |
Overall, GPT-4o shows relatively stable balanced accuracy throughout the phases, whereas the HINT model exhibits a progressively increasing trend in specificity. The applicability of different models varies at each phase; the HINT model performs more balancedly in the later stages, while GPT-4o is suitable for scenarios where high specificity is not a primary requirement due to its stable balanced accuracy.
@@ 2.3 Prediction on Clinical Trials Across Different Disease Categories
The distribution of diseases across different model datasets is shown in the figure. (Fig.1 A) Neoplasms consistently make up the largest portion of the dataset(30%-37%), while respiratory diseases have the smallest representation(5%-9%). Other categories, such as digestive system diseases, nervous system diseases, and endocrine/metabolic diseases, are also present in notable proportions across all models. Additionally, the ”Other” category (for example: Diseases of the eye and adnexa, Diseases of the ear and mastoid process, Diseases of the skin and subcutaneous tissue) appears prominently in the overall distribution. The models exhibit distinct strengths and limitations when predicting clinical trials across different disease categories (Table 3).For example, the HINT model achieved high average accuracy in ”Diseases of the respiratory system” and ”Diseases of the digestive system,” reaching about 0.75, while the large language models showed lower accuracy for these two categories, around 0.5.
The HINT model, while not leading in any specific category, showed stable performance, achieving a balanced accuracy of 0.704 and an MCC of 0.276 in digestive system diseases. A notable observation is the significant representation of oncology trials in the dataset, where none of the models performed well; all scored below their overall averages in both balanced accuracy and MCC. The HINT model, in particular, struggled in oncology trials with a low balanced accuracy (0.326) and MCC (0.141), suggesting that the high complexity and challenging nature of these trials negatively impacted model performance.
@@
In terms of recall, large language models consistently achieved high rates, indicating strong capability in identifying positive cases across various disease categories. However, the HINT model showed a low recall rate of 0.227 in oncology-related diseases, reflecting its limited ability to recognize positive cases in this challenging category. Conversely, for diseases such as endocrine, nutritional, and metabolic disorders, the HINT model achieved a higher recall rate of 0.875, indicating a stronger capacity for identifying positive cases within these categories.
@@
Despite the limited number of experiments, GPT-4o displayed strong recognition of negative samples, especially in respiratory and digestive diseases, where it identified all negative cases. Other models, like Llama, performed considerably worse, showing almost no detection of negative samples. Although the HINT model’s specificity was not as high as GPT-4o’s, it remained stable in respiratory and non-pathological diseases.
Overall, the HINT model showed consistent performance across various datasets. GPT-4mini and GPT-3.5, however, performed poorly across all disease categories, suggesting a lack of discriminatory ability. Similarly, while llama demonstrated some classification ability, it was limited and only evident in a few categories. The large proportion of oncology trials, coupled with the models’ limited performance in this category, likely contributed to their overall performance limitations. This analysis highlights that each model has particular strengths in specific disease categories, but the challenges posed by oncology trials significantly impact their overall effectiveness.
| Model | Disease Category | Trails Count | Balanced_acc | MCC | Recall | Specificity |
| Model: gpt4mini | ||||||
| Diseases of the respiratory system | 3 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Endocrine, nutritional and metabolic diseases | 6 | 1.000 | 0.0 | 1.000 | 0.0 | |
| Neoplasms | 18 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Diseases of the digestive system | 3 | 0.000 | 0.0 | 0.000 | 0.0 | |
| Diseases of the nervous system | 5 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Non-pathological diseases | 6 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Other | 2 | 0.500 | 0.0 | 1.000 | 0.0 | |
| Model: gpt4o | ||||||
| Diseases of the respiratory system | 3 | 1.000 | 1.000 | 1.000 | 1.000 | |
| Endocrine, nutritional and metabolic diseases | 6 | 0.833 | 0.000 | 0.833 | 0.000 | |
| Neoplasms | 18 | 0.526 | 0.081 | 0.909 | 0.143 | |
| Diseases of the digestive system | 3 | 0.333 | 0.000 | 0.000 | 0.333 | |
| Diseases of the nervous system | 5 | 0.667 | 0.408 | 1.000 | 0.333 | |
| Non-pathological diseases | 6 | 1.000 | 1.000 | 1.000 | 1.000 | |
| Other | 2 | 0.750 | 0.638 | 1.000 | 0.500 | |
| Model: gpt4 | ||||||
| Diseases of the respiratory system | 3 | 1.000 | 1.000 | 1.000 | 1.000 | |
| Endocrine, nutritional and metabolic diseases | 5 | 1.000 | 0.00 | 1.000 | 0.000 | |
| Neoplasms | 15 | 0.500 | 0.00 | 1.000 | 0.000 | |
| Diseases of the digestive system | 3 | 0.333 | 0.000 | 0.000 | 0.333 | |
| Diseases of the nervous system | 5 | 0.667 | 0.408 | 1.000 | 0.333 | |
| Non-pathological diseases | 5 | 0.500 | 0.000 | 1.000 | 0.000 | |
| Other | 2 | 0.786 | 0.670 | 1.000 | 0.571 | |
| Model: gpt3.5 | ||||||
| Diseases of the respiratory system | 148 | 0.491 | -0.060 | 0.983 | 0.000 | |
| Endocrine, nutritional and metabolic diseases | 173 | 0.500 | 0.000 | 1.000 | 0.000 | |
| Neoplasms | 712 | 0.501 | 0.012 | 0.996 | 0.005 | |
| Diseases of the digestive system | 192 | 0.525 | 0.155 | 0.994 | 0.056 | |
| Diseases of the nervous system | 241 | 0.497 | -0.040 | 0.994 | 0.000 | |
| Non-pathological diseases | 410 | 0.497 | -0.038 | 0.992 | 0.001 | |
| Other | 287 | 0.506 | 0.057 | 0.996 | 0.015 | |
| Model: HINT | ||||||
| Diseases of the respiratory system | 15 | 0.808 | 0.419 | 0.615 | 1.000 | |
| Endocrine, nutritional and metabolic diseases | 27 | 0.438 | -0.125 | 0.875 | 0.000 | |
| Neoplasms | 132 | 0.375 | -0.255 | 0.227 | 0.523 | |
| Diseases of the digestive system | 27 | 0.741 | 0.381 | 0.682 | 0.800 | |
| Diseases of the nervous system | 38 | 0.681 | 0.356 | 0.696 | 0.667 | |
| Non-pathological diseases | 74 | 0.576 | 0.130 | 0.595 | 0.556 | |
| Other | 70 | 0.575 | 0.117 | 0.718 | 0.431 | |
| Model: Llama | ||||||
| Diseases of the respiratory system | 23 | 0.562 | 0.181 | 0.923 | 0.200 | |
| Endocrine, nutritional and metabolic diseases | 35 | 0.484 | -0.053 | 0.969 | 0.000 | |
| Neoplasms | 124 | 0.531 | 0.116 | 0.954 | 0.108 | |
| Diseases of the digestive system | 27 | 0.496 | -0.009 | 0.850 | 0.143 | |
| Diseases of the nervous system | 39 | 0.492 | -0.024 | 0.893 | 0.091 | |
| Non-pathological diseases | 46 | 0.562 | 0.302 | 1.000 | 0.125 | |
| Other | 324 | 0.533 | 0.108 | 0.947 | 0.119 | |

### v2

@@ Abstract
This study evaluates the performance of large language models (LLMs) and the HINT model in predicting clinical trial outcomes, focusing on metrics including Balanced Accuracy, Matthews Correlation Coefficient (MCC), Recall, and Specificity. Results show that GPT-4o achieves superior overall performance among LLMs but, like its counterparts (GPT-3.5, GPT-4mini, Llama3), struggles with identifying negative outcomes. In contrast, HINT excels in negative sample recognition and demonstrates resilience to external factors (e.g., recruitment challenges) but underperforms in oncology trials, a major dataset component. LLMs exhibit strengths in early-phase trials and simpler endpoints like Overall Survival (OS), while HINT shows consistency across trial phases and excels in complex endpoints (e.g., Objective Response Rate). Trial duration analysis reveals improved model performance for medium-to-long-term trials, with GPT-4o and HINT displaying stability and enhanced specificity, respectively. We underscores the complementary potential of LLMs (e.g., GPT-4o, Llama3) and HINT, advocating for hybrid approaches to leverage GPT-4o’s predictive power and HINT’s specificity in clinical trial outcome forecasting.
@@ 2.1 Prediction on the Entire Clinical Trial Dataset
The test results across all models on the complete dataset are shown in the Table 1. We observed that the GPT-4o model performed the best in terms of balanced accuracy and MCC, reaching a balanced accuracy of 0.573 and an MCC of 0.212(Table 1). Although its recall was high at 0.931, its specificity remained relatively low at 0.214, suggesting a tendency to over-classify trials as successful while overlooking some potential failed trials.
GPT-4mini and GPT-4, while achieving perfect recall (1), exhibited extremely low specificity (0 and 0.059, respectively), indicating a strong bias toward predicting all cases as positive, resulting in balanced accuracies of 0.500 and 0.529. GPT-3.5 also demonstrated a similar trend with a recall of 0.998 and a specificity of 0.008, leading to a balanced accuracy of 0.503. This likely suggests that these three models have weak predictive power for the trial results across the overall dataset
The Llama3 model achieved moderate recall (0.964) but low specificity (0.092), resulting in a balanced accuracy of 0.528 and a low MCC of 0.116, indicating limited effectiveness in correctly identifying negative samples. However, it outperformed GPT-4, GPT-4mini, and GPT-3.5
Among all models, the HINT model achieved the highest specificity at 0.473, which contributed to its relatively strong balanced accuracy of 0.525 and an MCC of 0.050. Despite having lower recall and balanced accuracy compared to other models, HINT’s ability to identify negative cases suggests a more balanced performance across positive and negative samples.
@@
GPT-4mini | 0.500 | 0.000 | 1.000 | 0.000 |
GPT-4o | 0.532 | 0.098 | 0.913 | 0.150 |
GPT-4 | 0.529 | 0.171 | 1.000 | 0.059 |
GPT-3.5 | 0.503 | 0.046 | 0.998 | 0.008 |
HINT | 0.525 | 0.050 | 0.577 | 0.473 |
Llama3 | 0.528 | 0.116 | 0.964 | 0.092 |
@@ 2.2 Prediction of Different Clinical Trial Phases
As clinical trials progress from Phase I to Phase III, different models exhibit varying trends in metrics such as balanced accuracy, Matthews correlation coefficient (MCC), and specificity. For example, HINT’s metrics improve as the trial phase progresses, while Llama3 demonstrates stronger performance in predicting earlier trial phases.
In Phase I (Table 2), GPT-4o and GPT-4 achieve relatively high balanced accuracies of 0.557 and 0.600, respectively, but their specificities are low, at 0.40 and 0.20. This indicates certain limitations in identifying negative samples. Llama3 also demonstrates a relatively strong balanced accuracy of 0.556, though accompanied by a lower specificity. GPT-4mini shows poor performance across various metrics, indicating a lack of predictive power for clinical trials. HINT achieves a balanced accuracy of 0.470 and an MCC of -0.058 in Phase I, while GPT-3.5 scores 0.497 and -0.053, suggesting that both models have predictive capabilities weaker than random guessing at this stage. However, it is worth noting that HINT maintains the highest specificity of 0.412, indicating its continued strong ability to identify positive samples at this point.
In Phase II, the performance of GPT-4o declines, with its balanced accuracy dropping to 0.533, specificity decreasing to 0.067, although MCC shows a slight increase. The balanced accuracy of GPT-4 drops to 0.54, while its MCC decreases to 0.199 and specificity further decreases to 0.008. This downward trend is also observed in the Llama3 model. On the other hand, the HINT model’s balanced accuracy increases to 0.517, with improvements in all other metrics. Compared to Phase I, HINT shows improved predictive ability in Phase II clinical trials. GPT-3.5 also shows a slight improvement but does not demonstrate strong predictive power for clinical trials.
By Phase III, GPT-4o achieves its best performance, with balanced accuracy increasing to 0.625 and MCC reaching 0.408. Its specificity also improves to 0.25, continuing to display relatively stable performance. The HINT model further enhances its performance in this phase, achieving a balanced accuracy of 0.593 and an increase in specificity to 0.473. This demonstrates a balanced capability in recognizing both positive and negative samples. GPT-3.5 and Llama3 exhibit low specificities and weak abilities in identifying negative samples, while GPT-4mini’s specificity remains at 0 across all phases.
@@
1 | GPT-4mini | 0.500 | 0.000 | 1.000 | 0.000 |
2 | GPT-4o | 0.557 | 0.120 | 0.714 | 0.400 |
3 | GPT-4 | 0.600 | 0.333 | 1.000 | 0.200 |
4 | GPT-3.5 | 0.497 | -0.053 | 0.995 | 0.000 |
5 | HINT | 0.470 | -0.058 | 0.528 | 0.412 |
6 | Llama3 | 0.556 | 0.185 | 0.946 | 0.167 |
@@
1 | GPT-4mini | 0.500 | 0.000 | 1.000 | 0.000 |
2 | GPT-4o | 0.533 | 0.175 | 1.000 | 0.067 |
3 | GPT-4 | 0.542 | 0.199 | 1.000 | 0.083 |
4 | GPT-3.5 | 0.504 | 0.051 | 0.998 | 0.009 |
5 | HINT | 0.518 | 0.036 | 0.508 | 0.528 |
6 | Llama3 | 0.537 | 0.146 | 0.965 | 0.110 |
@@
1 | GPT-4mini | 0.500 | 0.000 | 1.000 | 0.000 |
2 | GPT-4o | 0.625 | 0.408 | 1.000 | 0.250 |
3 | GPT-4 | 0.500 | 0.000 | 1.000 | 0.000 |
4 | GPT-3.5 | 0.505 | 0.060 | 0.998 | 0.012 |
5 | HINT | 0.594 | 0.193 | 0.714 | 0.474 |
6 | Llama3 | 0.518 | 0.084 | 0.973 | 0.062 |
@@ 2.3 Prediction on Clinical Trials Across Different Disease Categories
The distribution of diseases across different model datasets is shown in Fig.1 A. Neoplasms consistently comprise the largest portion of the dataset(30%-37%), while respiratory diseases have the smallest representation(5%-9%). Other categories, such as digestive system diseases, nervous system diseases, and endocrine/metabolic diseases, are also present in notable proportions across all models. Additionally, the ”Other” category (for example: Diseases of the eye and adnexa, Diseases of the ear and mastoid process, Diseases of the skin and subcutaneous tissue) appears prominently in the overall distribution. The models exhibit distinct strengths and limitations when predicting clinical trials across different disease categories.For example, the HINT model achieved high average accuracy in ”Diseases of the respiratory system” and ”Diseases of the digestive system,” reaching about 0.75, while the large language models showed lower accuracy for these two categories, around 0.5.
The HINT model shows significantly lower performance in oncology trials, with a low balanced accuracy of 0.437 and an MCC of -0.137, suggesting that the high complexity and challenging nature of these trials negatively impacted the model’s performance. In contrast to previous observations, large language models did not exhibit such a disparity. Most large language models showed predictive abilities in oncology trials that were similar to their overall average performance.
@@
In terms of recall, large language models consistently achieved high rates, indicating a strong capability in identifying positive cases across various disease categories. However, the HINT model showed a low recall rate of 0.246 in oncology-related diseases, reflecting its limited ability to recognize positive cases in this challenging category. Conversely, the situation was the opposite in other disease categories; for diseases such as endocrine, nutritional, and metabolic disorders, the HINT model achieved a perfect recall rate of 1, indicating a stronger capacity for identifying positive cases within these categories.
@@
Despite the limited number of experiments, GPT-4o displayed strong recognition of negative samples, especially in respiratory and digestive diseases, where it identified many negative cases, achieving specific rates of 1 and 0.5, far higher than other categories. The Llama3 model, on the other hand, showed poor predictive performance in clinical trials for endocrine, nutritional, and metabolic diseases, as well as diseases of the digestive system, with balanced accuracy lower than random chance and almost no ability to identify negative samples in these categories.However, in other disease classifications, Llama3 has shown relatively excellent predictive ability, especially noteworthy is that: Llama3 demonstrated the strongest predictive ability for Neoplasms among all models, with an average accuracy of 0.555 and an MCC of 0.233. Other models, like GPT-3.5 and GPT-4mini, performed considerably worse, showing almost no detection of negative samples. Although the HINT model’s specificity was not as high as GPT-4o’s, it remained stable in respiratory and non-pathological diseases
Disease Category | Trails Count | Balanced_acc | MCC | Recall | Specificity |
Model: GPT-4mini | |||||
Diseases of the respiratory system | 3 | 0.500 | 0.0 | 1.000 | 0.0 |
Endocrine, nutritional and metabolic diseases | 6 | 0.500 | 0.0 | 1.000 | 0.0 |
Neoplasms | 18 | 0.500 | 0.0 | 1.000 | 0.0 |
Diseases of the digestive system | 3 | 0.500 | 0.0 | 0.000 | 0.0 |
Diseases of the nervous system | 5 | 0.500 | 0.0 | 1.000 | 0.0 |
Non-pathological diseases | 6 | 0.500 | 0.0 | 1.000 | 0.0 |
Other | 2 | 0.500 | 0.0 | 1.000 | 0.0 |
Model: GPT-4o | |||||
Diseases of the respiratory system | 3 | 0.750 | 0.500 | 0.500 | 1.000 |
Endocrine, nutritional and metabolic diseases | 6 | 0.250 | -0.632 | 0.500 | 0.000 |
Neoplasms | 18 | 0.512 | 0.040 | 0.900 | 0.125 |
Diseases of the digestive system | 3 | 0.750 | 0.500 | 1.000 | 0.500 |
Diseases of the nervous system | 5 | 0.200 | 0.000 | 0.000 | 0.200 |
Non-pathological diseases | 6 | 1.000 | 1.000 | 1.000 | 1.000 |
Other | 2 | 0.750 | 0.638 | 1.000 | 0.500 |
Model: GPT-4 | |||||
Diseases of the respiratory system | 3 | 0.750 | 0.500 | 0.500 | 1.000 |
Endocrine, nutritional and metabolic diseases | 5 | 0.500 | 0.00 | 1.000 | 0.000 |
Neoplasms | 15 | 0.500 | 0.00 | 1.000 | 0.000 |
Diseases of the digestive system | 3 | 0.750 | 0.500 | 0.100 | 0.500 |
Diseases of the nervous system | 5 | 0.200 | 0.000 | 0.000 | 0.200 |
Non-pathological diseases | 5 | 0.500 | 0.000 | 1.000 | 0.000 |
Other | 2 | 0.786 | 0.670 | 1.000 | 0.571 |
Model: GPT-3.5 | |||||
Diseases of the respiratory system | 148 | 0.501 | 0.011 | 0.988 | 0.015 |
Endocrine, nutritional and metabolic diseases | 173 | 0.500 | 0.000 | 1.000 | 0.000 |
Neoplasms | 712 | 0.504 | 0.064 | 1.000 | 0.008 |
Diseases of the digestive system | 192 | 0.525 | 0.155 | 0.994 | 0.056 |
Diseases of the nervous system | 241 | 0.504 | 0.061 | 1.000 | 0.008 |
Non-pathological diseases | 410 | 0.497 | -0.038 | 0.992 | 0.001 |
Other | 287 | 0.506 | 0.057 | 0.996 | 0.015 |
Model: HINT | |||||
Diseases of the respiratory system | 15 | 0.732 | 0.464 | 0.750 | 0.714 |
Endocrine, nutritional and metabolic diseases | 27 | 0.636 | 0.426 | 1.000 | 0.273 |
Neoplasms | 132 | 0.437 | -0.137 | 0.246 | 0.627 |
Diseases of the digestive system | 27 | 0.670 | 0.346 | 0.769 | 0.571 |
Diseases of the nervous system | 38 | 0.550 | 0.100 | 0.600 | 0.500 |
Non-pathological diseases | 74 | 0.576 | 0.130 | 0.595 | 0.556 |
Other | 70 | 0.575 | 0.117 | 0.718 | 0.431 |
Model: Llama3 | |||||
Diseases of the respiratory system | 23 | 0.575 | 0.219 | 0.929 | 0.222 |
Endocrine, nutritional and metabolic diseases | 35 | 0.479 | -0.116 | 0.958 | 0.000 |
Neoplasms | 124 | 0.555 | 0.223 | 0.985 | 0.125 |
Diseases of the digestive system | 27 | 0.495 | -0.015 | 0.846 | 0.143 |
Diseases of the nervous system | 39 | 0.554 | 0.178 | 0.950 | 0.158 |
Non-pathological diseases | 46 | 0.562 | 0.302 | 1.000 | 0.125 |
Other | 324 | 0.533 | 0.108 | 0.947 | 0.119 |